### v1

# Application of Battery Storage to Switching Predictive Control of Power Distribution Systems Including Road Heating*

Abstract
A road heating system is an electrical device which promotes snow melting by burying a heating cable as a thermal source underground. When integrating road heating into the power distribution system, we need to optimize the flow of electric power by appropriately integrating distributed power sources and conventional power distribution equipment. In this paper, we extend the power distribution system considered in the authors’ previous study to the case where battery storage is installed. As a main result, we propose a predictive switching control that achieves the reduction of distribution loss, attenuation of voltage fluctuation, and efficient snow melting, simultaneously. We verify the effectiveness of the application of battery storage through numerical simulation.
I Introduction
In heavy snowfall regions, there has long been a serious problem of great inconvenience and danger to people’s lives, for example falling on sidewalks, heavy workload due to snow removal, traffic congestion and etc [1][2]. A road heating system is one of the electrical devices which promotes snow melting by burying a thermal source underground, and raises the temperature of the ground surface [3]. However, the road heating system consumes a large amount of electrical power when we integrate road heating into the power distribution system [4][5]. We need to optimize the flow of electrical power by properly integrating distributed power sources, e.g. photovoltaic power generation (PV), and electric vehicles (EVs) with conventional power distribution equipment for an appropriate operation of power distribution systems.
Against the problem explained in the above paragraph, the authors of this paper [6] have derived a mathematical model of the power distribution system including road heating, and of snowmelting on the ground surface via thermal diffusion in the underground based on the nonlinear ordinary differential equation (ODE) model [7][8]. Furthermore, the authors proposed a predictive control by switching of the power distribution system including road heating [9]. The proposed control of  [9] achieves reduction of power distribution losses, attenuation of voltage impact, and efficient snowmelting, simultaneously. However, the supply of power electricity from the PV power generation equipment is limited to the distribution system only. There is a potential for improvement in terms of power flow optimization.
Based on the considerations given in the above paragraph, the purpose of this paper is to extend the results of the previous study [9] in the following two points.
  • We introduce a switching to the PV power generation equipment to achieve a more efficient power distribution system operation which does not depend on transmission from the grid.
  • A battery storage is introduced to the considered power distribution system to make effective use of surplus power generated by PV in the past.
Based on the above extension, this paper improves the optimization of power flow by applying the battery storage to the power distribution system, making more efficient use of the available power from the PV power generation equipment and the battery storage. As a main result, we give a design of predictive switching control of the distribution system under study, as well as a design of predictive control of the PV power generation equipment and battery storage. Finally, we verify the effectiveness of these predictive controls through numerical simulations. Through this verification, we show that the predictive control of the distribution system is more efficient than the previous study  [9].
The outline of this paper is described as follows. In Section II, we introduce a mathematical model of a power distribution system including road heating. As a main result, we give a switching predictive control for the system based on an evaluation function to simultaneously achieve reduction of distribution losses, attenuation of voltage impact, and efficient snowmelting in Section III. Finally, we verify an effectiveness of the proposed switching predictive control via numerical simulation in Section IV.
II Modeling of Power Distribution System
In this section, we develop the distribution system including road heating considered in a previous study due to the authors [6][9] to a mathematical model for the case including switching of the discharging of PV power generation equipment and battery storage.
II-A Overall setup of system
In this subsection, we describe the overall setup of the power distribution system including road heating. We show an overview of the system in Fig. 1. In the system, electrical power is supplied from the underground power distribution lines to the load electrical equipment (residences, EVs, lamps, and etc.) and from the PV electric power generation equipment to the underground power distribution lines. Assuming a realistic burial location, a heating cable is buried in parallel with the underground power distribution line in the horizontal direction. Two switches are connected at both terminals to control the voltage distribution over the heating cable. When Switch 1 is turned On, the power electricity is transmitted from the starting end of the heating cable. On the other hand, the power is transmitted from the battery storage if Switch 2 is turned On. In addition, the heat generated in the cable of the distribution line is transferred within the cable, and diffused into the underground. The Joule heat propagating to the ground surface is used to melt snow.
Figure 1: Considered power distribution system
We show a single-line diagram of the power distribution system in Fig. 2. The branches of the underground power distribution line and the heating cable at the starting point and terminal are denoted as Bifurcation points 1 and 2, respectively. We assume that Transformers A and B are connected to Switches 1 and 2, respectively.
Figure 2: Single-line diagram of the considered power distribution system
II-B Voltage distribution on underground power distribution line [7][8]
We consider the three-phase underground power distribution line where the line extends in a straight line from the transformer to the load electrical equipment and PV electric generation equipment between both bifurcation points as shown in Fig. 2. We also suppose that there are no voltage regulation devices such as transformers on the intermediate positions over the lines. The horizontal spatial variable on the underground power distribution line is defined as $x$ [m]. The starting point and terminal of the line correspond to $x=0$ [m] and $x=L_{\rm f}$ [m], $L_{\rm f}>0$, respectively.
II-B1 Voltage distribution on underground power distribution line
We describe the voltage distribution of an underground power distribution line, which is assumed to be in three-phase equilibrium, by the ordinary differential equations with position $x$ [m] as an independent variable, for an arbitrarily fixed time $t$ [min]. We denote the voltage phasor for single-phase at position $x$ and time $t$ by $v_{\rm e}(x;t)e^{j\theta_{\rm e}(x;t)}$, where $j$ is the imaginary unit. Then, $\theta_{\rm e}(x;t)$ [rad] and $v_{\rm e}(x;t)$ [V] correspond to the voltage phase and amplitude, respectively. The spatial variation on the distribution voltage profile is described by the nonlinear ODE model [7][8]:
$\frac{\mathrm{d}\theta_{\rm e}(x;t)}{\mathrm{d}x}$ $=-\frac{s_{\rm e}(x;t)}{v_{\rm e}(x;t)^{2}},$
$\frac{\mathrm{d}v_{\rm e}(x;t)}{\mathrm{d}x}$ $=w_{\rm e}(x;t),$
$\frac{\mathrm{d}s_{\rm e}(x;t)}{\mathrm{d}x}$ $=\frac{b_{\rm e}p_{\rm e}(x,t)-g_{\rm e}q_{\rm e}(x,t)}{g_{\rm e}^{2}+b_{\rm e}^{2}},$ (1)
$\frac{\mathrm{d}w_{\rm e}(x;t)}{\mathrm{d}x}$ $=\frac{s_{\rm e}(x;t)^{2}}{v_{\rm e}(x;t)^{3}}-\frac{g_{\rm e}p_{\rm e}(x,t)+b_{\rm e}q_{\rm e}(x,t)}{(g_{\rm e}^{2}+b_{\rm e}^{2})v_{\rm e}(x;t)},$ (2)
where $s_{\rm e}(x,t)$ [V2/m] and $w_{\rm e}(x,t)$ [V/m] represent the supplemental variable and voltage gradient at position $x$ and time $t$, respectively. In (1) and (2), $p_{\rm e}(x,t)$ [W/m] and $q_{\rm e}(x,t)$ [Var/m] are active and reactive powers at position $x$ and time $t$ of the underground power distribution line, respectively. If $p_{\rm e}(x,t)>0$ holds, $p_{\rm e}(x,t)$ denotes the active power, i.e. the supply, flowing to the distribution line at $x$. On the other hand, if $p_{\rm e}(x,t)<0$ holds, $p_{\rm e}(x,t)$ corresponds to the active power consumption flowing from the line at $x$. The same inequality holds for the reactive power $q_{\rm e}(x,t)$. Due to the assumptions made at the terminal of the underground power distribution line, $p_{\rm e}(L_{\rm f},t)=0$ and $q_{\rm e}(L_{\rm f},t)=0$ hold. The constants $g_{\rm e}$ [S/m] and $b_{\rm e}$ [S/m] are the conductance and susceptance per unit length at position $x$ of the line. Since the distribution loss $\Gamma_{\rm e}(x;t)$ [W/m] of the underground distribution line is considered for three phases, the distribution loss for a single-phase is three times larger, and is given by
$\Gamma_{\rm e}(x;t)=3g_{\rm e}\left(w_{\rm e}(x;t)^{2}+\frac{s_{\rm e}(x;t)^{2}}{v_{\rm e}(x;t)^{2}}\right)$
from [12].
II-B2 Boundary condition
The voltage phase and amplitude satisfy the following boundary conditions at each point:
Bifurcation point 1: $\theta_{\rm e}(0,t)=\theta_{1},\ v_{\rm e}(0,t)={v_{1}},$ (3)
Bifurcation point 2: $\theta_{\rm e}(L_{\rm f},t)=\theta_{2}(t),\ v_{\rm e}(L_{\rm f},t)=v_{2}(t),$ (4)
In (3), $\theta_{1}$ and $v_{1}$ denote the voltage phase and amplitude in Bifurcation point 1, respectively, which are supposed to be constant. Moreover, $\theta_{2}(t)$ and $v_{2}(t)$ are the voltage phase and amplitude in Bifurcation point 2 at time $t$, respectively, in (4).
II-C Voltage and temperature distribution of heating cable [6]
We assume that the heating cable is single-phase two-wire system. The heating cable extends in a straight line from Transformer A to Transformer B with the switches connected at both ends. The horizontal spatial variable of the heating cable is $x$ [m] in common with underground power distribution line. The vertical spatial variable $y$ [cm] is the depth from the ground surface to the heating cable. This implies that $y=0$ [cm] and $y=D_{\rm f}$ [cm] correspond to the ground and the surface, respectively.
II-C1 Voltage distribution
We denote the voltage phasor for single-phase at position $x$ and time $t$ by $v_{\rm h}(x;t)e^{j\theta_{\rm h}(x;t)}$, where $\theta_{\rm h}(x;t)$ [rad] and $v_{\rm h}(x;t)$ [V] are the voltage phase and amplitude, respectively. Similarly to the underground power distribution line, the spatial variation of the voltage distribution over the heating cable, which is assumed to be single-phase two-wire system, is described by the nonlinear ODE model [6].
$\frac{\mathrm{d}\theta_{\rm h}(x,t)}{\mathrm{d}x}$ $=-\frac{s_{\rm h}(x,t)}{v_{\rm h}(x,t)^{2}},$
$\frac{\mathrm{d}v_{\rm h}(x,t)}{\mathrm{d}x}$ $=w_{\rm h}(x,t),$
$\frac{\mathrm{d}s_{\rm h}(x,t)}{\mathrm{d}x}$ $=\frac{b_{\rm h}p_{\rm h}(x,t)-g_{\rm h}q_{\rm h}(x,t)}{g_{\rm h}^{2}+b_{\rm h}^{2}},$ (5)
$\frac{\mathrm{d}w_{\rm h}(x,t)}{\mathrm{d}x}$ $=\frac{s_{\rm h}(x,t)^{2}}{v_{\rm h}(x,t)^{3}}-\frac{g_{\rm h}p_{\rm h}(x,t)+b_{\rm h}q_{\rm h}(x,t)}{(g_{\rm h}^{2}+b_{\rm h}^{2})v_{\rm h}(x,t)},$ (6)
where $s_{\rm h}(x,t)$ [V2/m] and $w_{\rm h}(x,t)$ [V/m] represent the supplemental variable and voltage gradient at position $x$ and time $t$, respectively. In (5) and (6), $p_{\rm h}(x,t)$ [W/m] and $q_{\rm h}(x,t)$ [Var/m] are active and reactive powers at position $x$ and time $t$ of the heating cable, respectively. These powers satisfy the same properties that described for the underground power distribution line in the previous section. Moreover, $g_{\rm h}$ [S/m] and $b_{\rm h}$ [S/m] are the conductance and susceptance per unit length at position $x$ of the cable, respectively.
II-C2 Setting of transformers
We define $\theta_{\circ}(t)$ [rad], $v_{\circ}(t)$ [V], $s_{\circ}(t)$ [V2], and $w_{\circ}(t)$ [V/m] as the voltage phase, voltage amplitude, supplemental variable, and voltage gradient of Transformer $\circ$ $(\circ=\mathrm{A},\mathrm{B})$. At Transformer A, $\theta_{\mathrm{A}}(t)=\theta_{\mathrm{A}}$ and $v_{\mathrm{A}}(t)=v_{\mathrm{A}}$ are supposed to be constant. The turn ratio corresponds to the transformation ratio of the voltage amplitudes. The following equalities hold at each transformer from [8]:
Transformer A: $\theta_{\mathrm{h}}(0,t)=\theta_{\mathrm{A}},\ v_{\mathrm{h}}(0,t)=\frac{1}{a_{\rm A}}v_{\mathrm{A}},$
$s_{\mathrm{h}}(0,t)=s_{\mathrm{A}}(t),\ w_{\mathrm{h}}(0,t)=a_{\rm A}w_{\mathrm{A}}(t).,$
Transformer B: $\theta_{\mathrm{h}}(L_{\mathrm{f}},t)=\theta_{\mathrm{B}}(t),\ v_{\mathrm{h}}(L_{\mathrm{f}},t)=\frac{1}{a_{\mathrm{B}}}v_{\mathrm{B}}(t),$
$s_{\mathrm{h}}(L_{\mathrm{f}},t)=s_{\mathrm{B}}(t),\ w_{\mathrm{h}}(L_{\mathrm{f}},t)=a_{\rm B}w_{\mathrm{B}}(t).$
We describe the boundary conditions of the voltage phase and amplitude, supplemental variable, and voltage gradient at positions $x=0$ and $x=L_{\rm f}$ over the heating cable. The boundary conditions are described as follows corresponding to the switching patterns:
Switches 1 and 2: Off:
$\theta_{\rm h}(0;t)=0,\ v_{\rm h}(0;t)=0,\ s_{\rm h}(0;t)=0,\ w_{\rm h}(0;t)={0},$
$\theta_{\rm h}(L_{\rm f};t)=0,\ v_{\rm h}(L_{\rm f};t)=0,\ s_{\rm h}(L_{\rm f};t)=0,\ w_{\rm h}(L_{\rm f};t)=0,$
Switch 1: On:
$\theta_{\rm h}(0;t)=\theta_{1},\ v_{\rm h}(0;t)=\frac{1}{{a_{\rm 1}}}v_{1},$
$s_{\rm h}(L_{\rm f};t)=0,\ w_{\rm h}(L_{\rm f};t)=0,$
Switch 2: On:
$s_{\rm h}(0;t)=0,\ w_{\rm h}(0;t)=0,$
$\theta_{\rm h}(L_{\rm f};t)=\theta_{1},\ v_{\rm h}(L_{\rm f};t)=\frac{1}{{a_{\rm 1}}}v_{1}.$
In the above conditions, $\theta_{1}$ [rad] and $v_{1}$ [V] denote the reference voltage phase and amplitude for the underground distribution line at $x=0$. Moreover, $a_{\rm 1}$ is the transformer ratio of the single phase transformer.
II-C3 Thermal distribution of heating cable surface
At the surface of the heating cable, the transient spatial and temporal variation of the temperature satisfies the following ODE:
$\frac{\mathrm{d}\delta_{\rm surf,diff}(t;x)}{\mathrm{d}t}=\frac{\Gamma_{\rm h}(x,t)\times 10^{-2}-q_{\rm r}-\gamma_{\rm cable}\delta_{\rm surf,diff}(t;x)}{C_{\rm cable}}$
[…]
$\Gamma_{\rm h}(t;x)$ $=g_{\rm h}\left(w_{\rm h}(t;x)^{2}+\frac{s_{\rm h}(t;x)^{2}}{v_{\rm h}(t;x)^{2}}\right).$ (7)
The cable surface temperature $\delta_{\rm surf}(t;x)$ [C] at position $x$ and time $t$ is given by $\delta_{\rm surf}(t;x)=\delta_{\rm surf,diff}(t;x)+\delta_{\rm soil}(D_{\rm f},t;x)$, where $\delta_{\rm soil}(D_{\rm f},t;x)$ [C] is the temperature of the soil at position $x$ and depth $y=D_{\rm f}$. Based on the surface temperature and contact heat transfer [13], the boundary condition at the surface is given by
$-\lambda_{\rm tran}\frac{\mathrm{d}\delta_{\rm soil}(D_{\rm f},t;x)}{\mathrm{d}y}=\beta_{\rm cable}(\delta_{\rm surf}(x,t)-\delta_{\rm soil}(D_{\rm f},t;x)),$
[…]
II-D Switches
Since the voltages at Transformer B and the terminal $(x=L_{\rm f})$ of the underground power distribution line are required to be equal, if one switch is On, the other switch is Off. Moreover, if the heating cable is not used, both switches are Off. We define $\sigma_{i}(t)\in\{0,1\}$ $(i=0,1,2)$ as the binary variable by
$\sigma_{0}(t)$ :=$:=\begin{cases}1&\mbox{(Switches~{}1 and 2: {\tt Off})}\\ 0&\mbox{(Otherwise), }\\ \end{cases}$
$\sigma_{1}(t)$ :=$:=\begin{cases}1&\mbox{(Switch~{}1: {\tt On}, Switch~{}2: {\tt Off})}\\ 0&\mbox{(Otherwise), }\\ \end{cases}$
$\sigma_{2}(t)$ :=$:=\begin{cases}1&\mbox{(Switch~{}1: {\tt Off}, Switch~{}2: {\tt On})}\\ 0&\mbox{(Otherwise)}.\\ \end{cases}$
These variables can be regarded as the control input to the power distribution system. We refer to the integer $i$ as the index of the control input $\sigma_{i}(t)$.
We assume that the PV power generation equipment can transmit the generated electrical power to either underground power distribution line or battery storage. We define the binary variable $\sigma_{{\rm pv}}(t)\in\{0,1\}$ related to the power transmission for the equipment by corresponding to the direction of the power transmission as follows:
$\sigma_{{\rm pv}}(t)$ :=$:=\begin{cases}1&\mbox{(Underground distribution line)}\\ 0&\mbox{(Battery storage)}.\\ \end{cases}$
In the battery storage, it is assumed that the active power is supplied only from the PV power generation equipment. We assume that a certain amount of power can be transmitted to either the heating cable or the underground power distribution line, respectively. We also allow for the possibility of not transmitting power to either of them. The binary variable $\sigma_{{\rm battery},i}(t)\in\{0,1\}$ $(i=0,1,2)$ for the transmission of power from the battery storage is defined by the following equality corresponding to the direction of the power transmissions as the control input to the power distribution system:
$\sigma_{{\rm battery},0}(t)$ :=$:=\begin{cases}1&\mbox{(No power transmission)}\\ 0&\mbox{(Otherwise), }\\ \end{cases}$
$\sigma_{{\rm battery},1}(t)$ :=$:=\begin{cases}1&\mbox{(Underground distribution line)}\\ 0&\mbox{(Otherwise), }\\ \end{cases}$
$\sigma_{{\rm battery},2}(t)$ :=$:=\begin{cases}1&\mbox{(Heating cable)}\\ 0&\mbox{(Otherwise). }\\ \end{cases}$
II-E Thermal Diffusion in Underground and Snow Volume [6]
II-E1 Thermal diffusion
Throughout this paper, we assume that there is no horizontal thermal diffusion, and that the diffusion occurs only in the vertical direction in the underground. We define the underground temperature of the soil at position $x$, depth $y$ and time $t$ by $\delta_{\rm soil}(y,t;x)$ [C]. The spatial and temporal variation of the temperature is described by the following standard thermal conduction equation:
$\frac{\partial\delta_{\rm soil}(y,t;x)}{\partial t}$ $=\alpha_{\rm soil}\frac{\partial^{2}\delta_{\rm soil}(y,t;x)}{\partial y^{2}},$
where $\alpha_{\rm soil}$ [m2/s] is thermal diffusivity of soil, respectively. The boundary condition at the ground surface is given by the thermal conduction equation with the thermal exchange from the snow accumulation [13]:
$-\lambda_{\rm tran}\frac{\mathrm{d}\delta_{\rm soil}(0,t;x)}{\mathrm{d}y}=\beta_{\rm ground}(\delta_{\rm soil}(0,t;x)-\delta_{\rm snow}),$
where $\beta_{\rm ground}~{}\mathrm{[W/m\cdot K]}$ is the contact heat transfer coefficient per unit length between the ground surface and accumulation of snow. Moreover, $\delta_{\rm snow}$ [C] is the constant snow temperature on the ground surface.
II-E2 Snow melting
We suppose that the snow melting occurs due to the radiant heat from the sunlight and the Joule heat of the heating cable for simplicity. The spatial and temporal variation of the snow volume $h_{\rm snow}(x,t)$ [mm] is described by by the following equation [14]:
$\frac{\mathrm{d}h_{\rm snow}(t;x)}{\mathrm{d}t}$ $=-\frac{a_{\rm snow}}{d_{\rm snow}}(\mu_{1}(x,t)+\mu_{2}(x,t))+f_{\rm snow}(t).$ (8)
where $h_{\rm snow}(x,t)$ [mm] is the snow volume and $f_{\rm snow}(t)$ [mm/min] is the snowfall per 1 minute. Moreover, $d_{\rm snow}$ [g/cm3] is the density of snow volume, and $a_{\rm snow}$ is the unit conversion factor from the unit W/m2 to the unit mm/min. Finally, $\mu_{1}(x,t)$ [W/m2] represents the snowmelting due to the sunlight at position $x$ and time $t$. The snowmelting due to the influence of the sunlight is given by $\mu_{1}(x,t)=\phi_{\rm r}(t)+\phi_{\rm s}(x,t)+\phi_{\rm l}(x,t)$, where $\phi_{\rm r}(t)$ [W/m2] is the net radiation flux at time $t$, which is assumed to be independent on position $x$. Moreover, $\phi_{\rm s}(x,t)$ [W/m2] and $\phi_{\rm l}(x,t)$ [W/m2] denote the sensible and latent heat fluxes at position $x$ and time $t$, respectively. In addition, $\mu_{2}(x,t)$ [W/m2] represents the snowmelting due to the Joule heat of the cable at position $x$ and time $t$. Finally, in (8), the snowmelting due to the Joule heat of the heating cable is expressed as $\mu_{2}(x,t)=\beta_{\rm ground}(\delta_{\rm soil}(0,t;x)-\delta_{\rm snow})$.
III Switching Predictive Control
In this section, we explain the predictive control by switches connected to the heating cables and the predictive control of the PV power generation equipment and the battery storage as the main results of this paper.
III-A Design of Switching Predictive Control
In this susection, we explain the overall process for the predictive switching control of the power distribution system, the PV power generation equipment, and predictive control of battery storage. The optimal control inputs are computed according to Steps 0-5 below. Note that Step 0 and Steps 1-4 correspond to the preliminary and during operation of the predictive control, respectively. We also add Steps 3 and 4 to the predictive switching control of the previous study [9] due to the authors of this paper.
  • Step 0: We define the minimum switching time between the switching and the power transmission from the battery storage as $T_{\rm mini}$ [min]. Moreover, we define the predictive horizon as $T_{\rm pred}$ [min]. We prepare the time series data at each time  $t=kT_{\rm mini}$ $(k=0,1,2,\cdots)$ for the residential load, PV power generation, solar radiation, snowfall, temperature, wind speed, and snow accumulation as inputs for the predictive simulation.
  • Step 1: We define $N:=\frac{T_{\rm pred}}{T_{\rm mini}}$ as the total number of times at which switching can occur. This number corresponds to the sum of all $3^{N}$ switching patterns, $2^{N}$ PV transmission patterns, and $3^{N}$ discharging patterns of the battery storage. For all possible $3^{N}+2^{N}+3^{N}$ patterns, we perform the simulations from $t=kT_{\rm mini}$ to $t+T_{\rm pred}=(k+N)T_{\rm mini}$ based on the mathematical model of the distribution system.
  • Step 2: Based on the simulation in Step 1, we compute the value of the evaluation function $J[k]$ for the possible $3^{N}$ switching patterns ${\bf\it\mathbf{\sigma}}_{\rm d}[k]$ at $t=kT_{\rm mini}$. We then compare the $3^{N}$ values of the functions. We choose the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}[k]$ that has the smallest value at $t=kT_{\rm mini}$. We explain the detailed procedure of the design of the optimal switching pattern in Section III-B.
  • Step 3: Based on the simulation and the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}^{*}[k]$ computed in Steps 1 and 2, respectively, we compute the evaluation function for the $2^{N}$ possible transmission patterns of the PV power generation equipment at $t=kT_{\rm mini}$. We compute the value of the evaluation function $J_{\rm pv}[k]$ for ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ of the transmission patterns of the PV power generation equipment at $t=kT_{\rm mini}$. Then, by comparing the $2^{N}$ values of the evaluation functions, we find the optimal transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ of the PV power generation equipment whose value is the minimum at $t=kT_{\rm mini}$. The detailed procedure of this step is explained in Section III-C.
  • Step 4: Based on the predictive simulation in Step 1, the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}^{*}[k]$ in Step 2, and the optimal transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ in Step 3, we compute the values of the evaluation function $J_{\rm battery}[k]$ for the $3^{N}$ battery storage transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]$ at $t=kT_{\rm mini}$ for the evaluation function $J_{\rm battery}[k]$. In this case, the $3^{N}$ values of the evaluation function are compared, and the battery transfer pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]$ with the minimum value at $t=kT_{\rm mini}$ is selected as the optimal power transfer pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}^{*}[k]$. We explain the detailed procedure of the design of the optimal switching pattern in Section III-D.
  • Step 5: For the optimal switching pattern and the optimal transmission pattern computed in the previous steps, the control inputs are continuously applied from $t=kT_{\rm mini}$ to $t+T_{\rm mini}=(k+1)T_{\rm mini}$. Then, we choose the switching pattern which has the smallest value of the evaluation function as the optimal switching pattern. Fig. 3 illustrates the overall flow of the switching and power transmission selection from Step 2 to Step 4.
Figure 3: Selection tree of switches
III-B Optimal Switching of Heating Cable
In this subsection, we give the evaluation function that determines the switching on the heating cable. We define the evaluation function $J[k]$ at $t=kT_{\rm mini}$ by
$J[k]:=$ $w_{\rm loss}P_{\rm loss}[k]+w_{\rm fluc}V_{\rm fluc}[k]+w_{\rm snow}S_{\rm snow}[k]$
$+w_{\rm cost}M_{\rm cost}[k].$
The exact definitions and physical meaning of each term are explained in the remainder of this section.
[…]
    $P_{\rm loss}{[k]}$ $:={\sum_{l=1}^{N}\int_{(k+l-1)T_{\rm mini}}^{(k+l)T_{\rm mini}}}\int_{0}^{L_{\rm f}}\Gamma(x;t)\mathrm{d}x\mathrm{d}t,$
    $\Gamma(x;t)$ $:=\Gamma_{\rm e}(x;t)+\Gamma_{\rm h}(x;t).$ (9)
    In (9), $\Gamma_{\rm e}(x;t)$ [W/m] and $\Gamma_{\rm h}(x;t)$ [W/m] represent the distribution loss of the underground power distribution line and heating cable, respectively. Since the distribution loss of underground distribution line is considered for three phases, the distribution loss for a single-phase is three times larger and is given by
    $\Gamma_{\rm e}(x;t)=3g_{\rm e}\left(w_{\rm e}(x;t)^{2}+\frac{s_{\rm e}(x;t)^{2}}{v_{\rm e}(x;t)^{2}}\right)$
    from [12], and $\Gamma_{\rm h}(x,t)$ [W/m] is given by (7).
  • The second term: The function $V_{\rm fluc}{[k]}$ [V] corresponds to the evaluation of the voltage fluctuation of the underground power distribution line in the time interval $[t,t+T_{\rm pred})$, and $w_{\rm fluc}$ is the weighting coefficient in the evaluation of the fluctuation. The function $V_{\rm fluc}[k]$ is defined by the maximum absolute values of the difference between the voltage amplitude and the reference value $v_{1}$ [V] for $N$ intervals:
[…]
    $S_{\rm snow}{[k]}:={\sum_{{l}=1}^{N}\int_{(k+l-1)T_{\rm mini}}^{(k+l)T_{\rm mini}}}\int_{0}^{L_{\rm f}}h_{\rm snow}(x,t)\mathrm{d}x\mathrm{d}t,$
    where $w_{\rm snow}$ is the weighting coefficient in the evaluation of the volume.
  • The fourth term: The function $M_{\rm cost}[k]$ corresponds to the cost of operating the heating cable during the time interval $[t,t+T_{\rm pred})$, and $w_{\rm cost}$ represents the weighting coefficient of its evaluation in the evaluation function. As we suppose that the owner of the heating cable, the PV power generation equipment, and the battery storage is supposed to be the same. the cost occurs only when Switch 1 is set to On. Thus, the cost $M_{\rm cost}[k]$ is defined as
    $M_{\rm cost}{[k]}:={\sum_{{l}=1}^{N}\int_{(k+l-1)T_{\rm mini}}^{(k+l)T_{\rm mini}}}m_{\rm cost}(t)\mathrm{d}t.$
    In the above equality, $m_{\rm cost}(t)$ is defined by
    $m_{\rm cost}(t):=\begin{cases}-p_{\rm h}(x_{\rm load},t)&\mbox{(Switch~{}1: {\tt On})}\\ 0&\mbox{(otherwise)},\\ \end{cases}$
    where $p_{\rm h}(x_{\rm load},t)$ [W/m], $p_{\rm h}(x_{\rm load},t)\leq 0$ is the active power of the loads over the heating cable. Moreover, $x_{\rm load}$ corresponds to one of connection positions of the load on the heating cable except for the starting point $x=0$ and the terminal $x=L_{\rm f}$.
We define the discrete-time vector ${\bm{\sigma}}_{\rm d}[k]$, which consists of switching patterns from $t=kT_{\rm mini}$ to $t+T_{\rm pred}=(k+N)T_{\rm mini}$, by
${\bm{\sigma}}_{\rm d}[k]:=\begin{bmatrix}{\bm{\sigma}}_{{\rm d},0}[k]\\ {\bm{\sigma}}_{{\rm d},1}[k]\\ {\bm{\sigma}}_{{\rm d},2}[k]\\ \end{bmatrix},\ {\bm{\sigma}}_{{\rm d},i}[k]:=\begin{bmatrix}\sigma_{{\rm d},i}[k]\\ \sigma_{{\rm d},i}[k+1]\\ \vdots\\ \sigma_{{\rm d},i}[k+N]\\ \end{bmatrix}\ (i=0,1,2).$
We enumerate all $3^{N}$ patterns ${\bf\it\mathbf{\sigma}}_{\rm d}[k]$, and perform a numerical simulation for each of them. We obtain the optimal switching pattern ${\bm{\sigma}}_{\rm d}^{*}[k]\in\left\{0,1\right\}^{N}$ at $t=kT_{\rm mini}$ by solving the binary minimization problem
${\bm{\sigma}}{{}^{*}[k]}=\underset{{\bm{\sigma}}[k]\in\left\{0,1\right\}^{N}}{\operatorname{argmin}}\ J[k].$
From the current time $t=kT_{\rm mini}$ to the next switching time $t+T_{\rm mini}=(k+1)T_{\rm mini}$, we continue to apply the following optimal control input $\sigma_{{\rm opt},i}(t)$ $(i=0,1,2)$ by
$\sigma_{{\rm opt},i}(t)$ $=\sigma_{{\rm d},i}^{*}(kT_{\rm mini}),\ t\in[kT_{\rm mini},(k+1)T_{\rm mini}).$
III-C Optimal Switching of PV Power Generation Equipment
In this subsection, we formulate the evaluation function that determines the destination of power electricity from the PV power generation equipment. Specifically, this function $J_{\rm pv}[k]$ is mathematically defined by
$J_{\rm pv}[k]:=w_{\rm pvfluc}V_{\rm pvfluc}[k]-w_{\rm stor,1}B_{\rm stor,1}[k].$
  • The first term: The function $V_{\rm pvfluc}[k]$ [V] corresponds to the evaluation of the voltage fluctuation of the underground power distribution line in the time interval $[t,t+T_{\rm pred})$, and $w_{\rm pvfluc}$ represents the weighting coefficient in the evaluation function. The function $V_{\rm pvfluc}[k]$ is defined in the same way as in the equation (10).
  • The second term: The function $B_{\rm stor,1}[k]$ corresponds to the evaluation of the battery power in the time interval $[t,t+T_{\rm pred})$, and $w_{\rm stor,1}$ is the weighting coefficient in the evaluation function. Since the owner of the PV power generation equipment and the battery storage are assumed to be the same, this function corresponds to smoothing for the amount of power electricity $P_{\rm battery}(t)$ [kWh] of the battery storage. Thus, we define $B_{\rm stor,1}[k]$ to be large when the amount of power electricity in the battery storage is small. On the other hand, we set $B_{\rm stor,1}[k]$ to be small when the amount of power electricity in the battery storage is large. In this case, $B_{\rm stor,1}{[k]}$ can be defined by the following equality:
    $B_{\rm stor,1}{[k]}:={\sum_{{l}=1}^{N}\int_{(k+l-1)T_{\rm mini}}^{(k+l)T_{\rm mini}}}\zeta_{\rm stor,1}(t)\mathrm{d}t,$
    where $\zeta_{\rm stor,1}(t)$ is defined as follows corresponding to the power transmissions:
[…]
We describe the design of the predictive control of PV power generation equipment. We define the vector consisting of the transmission patterns from time $t=kT_{\rm mini}$ to $t+T_{\rm pred}=k+NT_{\rm mini}$ by
${\bf\it{\bf\sigma}}_{{\rm d,pv}}[k]:=\begin{bmatrix}\sigma_{{\rm d,pv}}[k]\\ \sigma_{{\rm d,pv}}[k+1]\\ \vdots\\ \sigma_{{\rm d,pv}}[k+N]\\ \end{bmatrix}.$
We enumerate all $2^{N}$ possible transmission patterns ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$, and compute the value of the evaluation function $J_{\rm pv}[k]$ corresponding to each transmission pattern. Then, the optimal transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,pv}^{*}[k]\in\left\{0,1\right\}^{N}$ is expressed as
${\bf\it\mathbf{\sigma}}_{\rm d,pv}^{*}[k]=\underset{{\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]\in\left\{0,1\right\}^{N}}{\operatorname{argmin}}\ J_{\rm pv}[k].$
During the time interval from time $t=kT_{\rm mini}$ to the next switching time $t+T_{\rm mini}=(k+1)T_{\rm mini}$, the control input $\sigma_{{\rm opt,pv}}(t)$ is given by
$\sigma_{{\rm opt,pv}}(t)$ $=\sigma_{{\rm d,pv}}^{*}(kT_{\rm mini}),\ t\in[kT_{\rm mini},(k+1)T_{\rm mini}).$
III-D Optimal Discharging of Battery Storage
In this subsection, We provide an evaluation function that determines the destination of the power discharged from the battery storage. We define the evaluation function $J_{\rm battery}[k]$ by
$J_{\rm battery}[k]:=w_{\rm batteryfluc}V_{\rm batteryfluc}[k]+w_{\rm stor,2}B_{\rm stor,2}[k].$
  • The first term: The function $V_{\rm batteryfluc}[k]$ [V] correspond to the evaluation of the voltage fluctuation of the underground distribution line in the time interval $[t,t+T_{\rm pred})$, and is defined in the same way as in the definition (10). Moreover, $w_{\rm batteryfluc}$ is the weighting coefficient of its evaluation.
  • The second term: The function
    $B_{\rm stor,2}{[k]}:={\sum_{{l}=1}^{N}\int_{(k+l-1)T_{\rm mini}}^{(k+l)T_{\rm mini}}}\zeta_{\rm stor,2}(t)\mathrm{d}t$
    corresponds to the evaluation of the amount of electricity in battery storage in the time interval $[t,t+T_{\rm pred})$, where $\zeta_{\rm stor,2}(t)$ is defined as follows corresponding to the power transmissions:
[…]
    In addition, $w_{\rm stor,2}$ represents the weighting coefficient of the evaluation.
We define the vector consisting of the discharging pattern from $t=kT_{\rm mini}$ to $t=k+T_{\rm pred}=k+NT_{\rm min}$ by
${\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]$ $:=\begin{bmatrix}{\bf\it\mathbf{\sigma}}_{{\rm d,battery},0}[k]\\ {\bf\it\mathbf{\sigma}}_{{\rm d,battery},1}[k]\\ {\bf\it\mathbf{\sigma}}_{{\rm d,battery},2}[k]\\ \end{bmatrix},$
${\bf\it{\bf\sigma}}_{{\rm d,battery},i}[k]$ $:=\begin{bmatrix}\sigma_{{\rm d,battery},i}[k]\\ \sigma_{{\rm d,battery},i}[k+1]\\ \vdots\\ \sigma_{{\rm d,battery},i}[k+N]\\ \end{bmatrix}\ (i=0,1,2).$
We enumerate all $3^{N}$ patterns of ${\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]$, and perform predictive simulations of the switching control. Then, the optimal discharging pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}^{*}[k]\in\left\{0,1\right\}^{N}$ at $t=kT_{\rm mini}$ is expressed as
${\bf\it\mathbf{\sigma}}_{\rm d,battery}^{*}[k]=\underset{{\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]\in\left\{0,1\right\}^{N}}{\operatorname{argmin}}\ J_{\rm battery}[k].$
Then, during the time interval from $t=kT_{\rm mini}$ to the next switching time $t+T_{\rm mini}=(k+1)T_{\rm mini}$, we continue to apply the optimal control input
$\sigma_{{\rm opt,battery},i}(t)=\sigma_{{\rm d,battery},i}^{*}(kT_{\rm mini})\ (i=0,1,2)$ ,
$t\in[kT_{\rm mini},(k+1)T_{\rm mini}).$
IV Numerical Verification
In this section, we verify the effectiveness of the proposed predictive switching control via numerical simulations considering two cases of typical power consumption, PV electric power generation, and weather conditions. These simulations suppose the winter season of Toyama Prefecture which is one of the heaviest snowfall areas in Japan.
IV-A Simulation Setting
In this subsection, we provide the setting of the power distribution system considered in the numerical simulation. We consider a power distribution system consisting of the underground power distribution line and the heating cable with length $100$ m. The single-line diagram of the system is illustrated in Fig. 4. For the underground power distribution line, the reference voltage amplitude for p.u. value is set to $v_{1}=\frac{6600}{\sqrt{3}}$ [V]. Since we deal with only one of the three phases of the underground distribution line in this paper, we convert this reference value of line voltage  $\sqrt{3}v_{1}=6600$ [V] to a phase voltage. In the heating cable, the reference voltage amplitude for the p.u. value is set to $\frac{1}{a_{1}}v_{1}=200$ [V], where $a_{1}=\frac{33}{\sqrt{3}}$ is the transformer ratio. The load electrical equipment is installed at the horizontal position $x=25$ [m] from Bifurcation point 1, and the PV electric power generation equipment is installed at $x=50$ [m]. In addition, the battery storage is connected at $x=75$ [m]. The connecting position $x_{\rm load}$ [m] of the load of the heating cable is given by
$\begin{cases}\text{No load connected}&\text{(Switches~{}1 and 2: {\tt Off})}\\ x_{\rm load}\fallingdotseq L_{\rm f}&\text{(Switch~{}1: {\tt On})}\\ x_{\rm load}\fallingdotseq 0&\text{(Switch~{}2: {\tt On}). }\end{cases}$
Figure 4: Single-line diagram of power distribution system considered in numerical simulation
We describe the parameters used in the simulations of this section. The parameters related to the voltage distribution in the underground power distribution line and heating cable are shown in Tables I and II, respectively. For underground power distribution line and heating cable, the reference voltage amplitudes for the p.u. values are $v_{1}=6600$ [V] and $\frac{1}{a_{\rm A}}v_{1}=200$ [V], respectively. We assume that no reactive power is supplied at all position $x$ of the underground power distribution line and heating cable, i.e. $q_{\rm e}(x,t)=q_{\rm h}(x,t)=0$ [p.u.]. The parameters related to the temperature distribution of the heating cable, the thermal diffusion in the underground, and the snowmelting on the ground surface are shown in Table III. The parameters related to the evaluation function are summarized in Table IV. We set the prediction horizon to be $T_{\rm pred}=30$ [min] from the viewpoint of the prediction of temporal changes of snow melting against the occurrence of pedestrian falls and traffic congestion due to snow accumulation. Moreover, the minimum switching time is set to $T_{\rm mini}=10$ [min], considering the influence of frequent switching on the power distribution system.
TABLE I: Reference values of voltage profile [15]
Parameter Symbol Value
Apparent power $W_{\rm base}$ $10$ [kVA]
Length of underground
power distribution line
$L_{\rm f}$ $100$ [m]
Voltage phase of underground
power distribution line
$\theta_{1}$ 00 [rad]
Voltage amplitude of underground
power distribution line
$v_{1}$ $\frac{6600}{\sqrt{3}}$ [V]
Voltage amplitude
of heating cable
$\frac{1}{{a_{\rm 1}}}v_{1}$ $200$ [V]
Transformer ratio ${a_{\rm 1}}$ $\frac{33}{\sqrt{3}}$
Conductance $g_{\rm base}$ $918$ [m/$\Omega$]
Susceptance $b_{\rm base}$ $918$ [m/$\Omega$]
TABLE II: p.u. values of voltage profile [15]
Parameter Symbol p.u. value
Underground power distribution line
Length $L_{\rm f}$ $1$
Voltage amplitude $v_{1}$ $1$
Conductance $g_{\rm e}$ $1$
Susceptance $b_{\rm e}$ $1$
Heating cable
Length $L_{\rm f}$ $1$
Conductance $g_{\rm h}$ $0.5$
Susceptance $b_{\rm h}$ $0.5$
Active power of load $p_{\rm h}(x_{\rm load},t)$ $-10$
TABLE III: Parameters related to heating cable, thermal diffusion, and snowmelting [14]
Parameter Symbol Value
Heating cable
Radiative cooling $q_{\rm r}$ 00 [W/cm]
Contact heat transfer coefficient
per unit length
between cable surface and soil
[…]
Heat capacity $C_{\rm cable}$ $18$ [$\cdot{}^{\circ}$]
[…]
Heat transfer coefficient
between heating cable and soil
[…]
Thermal diffusion
Burial depth of heating cable $D_{\rm f}$ $10$ [cm]
Thermal diffusivity of soil $\alpha_{\rm soil}$ $0.008$ [m2/s]
Heat transfer coefficient
between ground surface
and snow
[…]
Snowmelting
Snow temperature $\delta_{\rm snow}$ 00 [C]
Initial condition for snow volume $h_{\rm snow}(x,0)$ $30$ [mm]
Unit conversion factor $a_{\rm snow}$ $1.792\times 10^{-4}$
Density of snow volume $d_{\rm snow}$ $0.06$ [g/cm3]
TABLE IV: Parameters related to evaluation function
Parameter Symbol Value
Minimum switching time $T_{\rm mini}$ $10$ [min]
Predictive horizon $T_{\rm pred}$ $30$ [min]
Weighting coefficients (Heating cable)
Distribution loss $w_{\rm loss}$ $4\times 10^{2}$
Voltage fluctuation $w_{\rm fluc}$ $1\times 10^{7}$
Snow volume $w_{\rm snow}$ $1.2\times 10^{5}$
Cost $w_{\rm cost}$ $8\times 10^{5}$
Weighting coefficients (PV)
Voltage fluctuation $w_{\rm pvfluc}$ $1$
Amount of power of battery storage $w_{\rm stor,1}$ $1\times 10^{-3}$
Weighting coefficients (Battery storage)
Voltage fluctuation $w_{\rm batteryfluc}$ $1$
Amount of power of battery storage $w_{\rm stor,2}$ $1\times 10^{-3}$
IV-B Time Series Data for Simulation
In this section, we describe the acquisition and processing of actual time series data on electricity and weather used in the simulations of this chapter.
IV-B1 Power consumption and PV electric power generation
The time series data on the power consumption of the residential loads are provided by Nagoya University Open Data for EMS Evaluation [16] for ten residential houses. As we consider only single phase of the three-phase underground distribution line, 1/3 of the time series data for 10 houses is used as the active power $p_{\rm e}(x,t)$ of the residential load.
We use time series data of PV electric power generation provided by Hokuriku Electric Power Transmission and Distribution Company [17]. We also use the value that is $\frac{1}{1000}$ of the data in order to fit our data to the scale of the underground power distribution line considered in this paper. Furthermore, as we consider only one phase among the three phases, we use 1/3 of the time series data that fits the scale of the distribution system considered in this paper as the effective power $p_{\rm e}(x,t)$ of the PV power generation equipment. On the other hand, when the power electricity is transmitted from the PV power generation equipment to the battery storage, we suppose that all the power electricity is transmitted. For this reason, we do not modify the scale of the data which is fit to the scale considered in this paper.
The power supplied by the PV electric power generation is considered to be more accurately predicted based on the day’s weather forecast, compared to the power consumption of the load. Thus, we use the time series data for PV electric power generation as predictive data. In general, it is not easy to predict these data exactly. For this reason, the predicted value of the data at the predictive horizon continues to use the value at the current time as a constant value in the computation of the evaluation function.
IV-B2 Weather
As the time series data of solar radiation, we use the average data of the radiation from 1981 to 2009 provided by NEDO’s database viewing system [18]. Since this data is time series data at 1 h intervals, it is linearly interpolated as data at 30 sec intervals. Moreover, we choose the time series data observed in Toyama City. The data is provided by the Japan Meteorological Agency [19] for atmospheric temperature and wind speed. Since this data is given as time series data every 10 min, linear interpolation is performed to obtain data every 30 sec.
IV-C Case 1 (Morning without snowfall)
IV-C1 Setting and time series data
In this subsection, we consider a numerical simulation for a sunny winter morning with no snowfall from A.M. 6:00 to 8:00. This implies that that A.M. 6:00 and 8:00 correspond to $t=0$ [min] and $t=120$ [min], respectively. For the time series data used in this case, we use January 19, 2016 for the residential load [16], January 19, 2020 for the PV electric power generation [17]. Fig. 5 shows the temporal variation of the power consumption of residential loads and PV electric power generation. We also use the time series data for atmospheric temperature [19], wind speed [19], and the average data from January 19, 1981 to 2009 for solar radiation [18]. The time series data on these quantities are shown in Fig. 6.
(a) Power consumption of residential loads
(b) PV electric power generation
Figure 5: Temporal variation of active powers (Case 1)
(a) Solar radiation
(b) Snowfall
(c) Atmospheric temperature
(d) Wind speed
Figure 6: Temporal variation of variables related to weather (Case 1)
IV-C2 Simulation result
Temporal variation of variables
We show the simulation results for Case 1 in Fig. 7. In Fig. 7 (a), the voltage amplitude $v_{\rm e}(x;t)$ [V] of the underground power distribution line varies as power consumption of the residential load and heating cable, and power supply from PV. The amplitude is reduced due to the power consumption of the residential load during A.M. 6:50 to 7:00. On the other hand, the amplitude increases with the PV electric power generation between A.M. 7:50 and 8:00. Fig. 7 (b) shows that the voltage amplitude $v_{\rm h}(x;t)$ [V] of the heating cable varies greatly as the switches change. When Switch 1 is turned On, the voltage amplitude decreases from $x=0$ [m] to $x=100$ [m]. On the other hand, the amplitude decreases from $x=100$ [m] to $x=0$ [m] during Switch 2 is turned On.
The temperature $\delta_{\rm surf}(t;x)$ [C] of the cable surface is shown in Fig. 7 (c). The temperature increases because Switches 1 or 2 are turned On from A.M. 6:00 to 7:30, and decreases during A.M. 7:30 to 8:00 when both switches are turned Off. When Switch 1 is turned On, the temperature at $x=0$ [m] is higher than at $x=100$ [m]. On the other hand, if Switch 2 is turned On, the temperature at $x=100$ [m] increases compared to the temperature at $x=0$ [m]. From Fig. 7 (d), the snow volume $h_{\rm snow}(t;x)$ [mm] decreases with time.
(a) Voltage amplitude of underground distribution line
(b) Voltage amplitude of heating cable
(c) Temperature of cable surface
(d) Snow volume
Figure 7: Spatial and temporal variation of the variables (Case 1)
Temporal variation of switching and discharging
We show the simulation results for switching, and the PV power generation equipment, battery storage in Fig. 8.
Fig. 8 (a) illustrates the temporal variation of the switching control input for Case 1 by showing the index $i$ which satisfies $\sigma_{i}(t)=1$, the switch that turns On. From A.M. 6:00 to A.M. 6:30, Switch 2, which is powered by a battery storage, is turned to On. It is assumed that the owners of the heating cable and battery storage are the same, and Switch 2 is selected because it is less expensive than Switch 1. From A.M. 6:30 to 7:20, Switch 1 is set to On to melt snow, because the selection of Switch 2 would violate the constraint on the reserve capacity of the battery storage. From A.M. 7:20 to 7:30, Switch 2 is used for snow melting because the PV power generation system has supplied power to the battery storage. From A.M. 7:30 a.m. to 8:00, the switch is set to Off because the snow melting is almost completed.
Next, we show the temporal variation of the control input $\sigma_{{\rm pv}}(t)$ of the PV power generation equipment in Fig. 8 (b). It can be confirmed that the power electricity is transmitted to the battery storage from A.M. 6:00 to 7:20 and from A.M. 7:40 to 7:50, and to the underground power distribution line from A.M. 7:20 to 7:40 and from A.M. 7:50 to 8:00.
Finally, we show the temporal variation of the index $i$ satisfying $\sigma_{{\rm battery},i}(t)=1$, i.e. the destination of the power electricity of the battery storage in Fig. 8 (c). When Switch 2 is turned On, the battery storage always transmits the power electricity to the heating cable. In addition, from A.M. 6:30 to 7:20 and from A.M. 7:30 to 8:00, the power electricity is stored to the battery without being transmitted anywhere. We also present the temporal variation of the amount of electricity in the battery storage in Fig. 8 (d). From A.M. 6:00 a.m. to 6:30 and from A.M. 7:20 a.m. to 7:30, the amount of power electricity decreases with time because the battery storage transmits power to the heating cables. From A.M. 7:10 to 7:20 a.m. and from A.M. 7:40 to 7:50 a.m., the amount of power electricity increases because the PV power generation equipment transmits power while the battery storage is not transmitting power electricity anywhere.
(a) Temporal variation of switching control input
(b) Temporal variation of the destination of PV power
(c) Temporal variation of the destination of battery power
(d) Temporal variation of the amount of electricity in battery
Figure 8: Temporal variation with respect to switches, PV, and battery storage (Case 1)
Comparison with and without battery storage
In the following, we compare the cost and voltage variations without and with battery storage. Note that the former and latter simulations correspond to the result of the switching control due to the authors’ previous study [9] and this paper, respectively.
[…]
Figs. 9 (c) and (d) show the temporal variation of the voltage fluctuation of the underground power distribution line without and with the battery storage, respectively. Comparing the temporal variations for each case, there is no significant difference between the two cases. In Fig. 9 (c), if no battery storage is installed, the PV power generation equipment transmits all the power electricity to the underground power distribution line. On the other hand, in Fig. 9 (d), the equipment transmits power to the underground power distribution line and the battery storage. This shows the fact that the installation of the storage efficiently suppresses the voltage fluctuations while recharging the power, which implies the efficiency of the switching control of this paper comparing with [9].
(a) Temporal variation of switching control input (without battery storage)
(b) Temporal variation of switching control input (with battery storage)
(c) Voltage fluctuation (without battery storage)
(d) Voltage fluctuation (with battery storage)
Figure 9: Comparison of switching control input and voltage fluctuation with respect to battery storage (Case 1)
IV-D Case 2 (Evening with snowfall)
IV-D1 Setting and time series data
In this subsection, we consider a numerical simulation for a winter evening from P.M. 4:00 to 6:00 with snowfall starting from P.M. 5:00. This implies that P.M. 4:00 and 6:00 correspond to $t=0$ [min] and $t=120$ [min], respectively. We use the time series data of the residential load [16] from February 7, 2016, and of the PV electric power generation [17] from February 7, 2020. The temporal variation of the power consumption of residential load and PV electric power generation is shown in Fig. 10. The time series data for atmospheric temperature and wind speed are from February 7, 2020 [19]. Moreover, we choose the average solar radiation data from February 7, 1981 to 2009 [18]. The snowfall time series data gives a hypothetical snowfall of $12$ mm per $1$ hour assumed for Toyama Prefecture. The time series data of physical quantities related to weather are shown in Fig. 11.
(a) Power consumption of residential loads
(b) PV electric power generation
Figure 10: Temporal variation of active powers (Case 2)
(a) Solar radiation
(b) Snowfall
(c) Atmospheric temperature
(d) Wind speed
Figure 11: Temporal variation of variables related to weather (Case 2)
IV-D2 Simulation result
Temporal variation of variables
The simulation results for Case 2 are shown in Fig. 12. We show the temporal and spatial variations of the voltage amplitude of the underground distribution line in Fig. 12 (a). In this figure, the voltage fluctuation is suppressed by the power supplied by the PV power generation equipment against the increase of power consumption by the residential load from P.M. 5:00 to 5:30. In Fig. 12 (b), the voltage amplitude of the heating cable decreases from $x=100$ [m] to $x=0$ [m] because Switch 2 is set to On at many time intervals.
We show the temporal and spatial variations of the outer surface temperature  $\delta_{\rm surf}(t;x)$ [C] of the heating cable in Fig. 12 (c). As for the temperature at $x=100$ [m], the temperature at $x=0$ [m] is higher than that at $x=0$ [m], because Switch 2 is set to On during many time periods. Moreover, in Fig. 12 (d), The occurrence of snowfall from P.M. 5:00 has changed the decrease of the snow accumulation $h_{\rm snow}(t;x)$ [mm].
(a) Voltage amplitude of underground distribution line
(b) Voltage amplitude of heating cable
(c) Temperature of cable surface
(d) Snow volume
Figure 12: Spatial and temporal variation of the variables (Case 2)
Temporal variation of switching and discharging
We show the simulation results for the switches, the PV power generation equipment, and the battery storage in Fig. 13. The temporal variation of the switching of the switch with subscript $i$, i.e., On, satisfying $\sigma_{{\rm b},i}(t)=1$ is shown in Fig. 13 (a). From P.M. 4:00 to 5:30, Switch 2, which is powered by a battery storage, is set to On. Since the owners of the heating cables and the battery storage are supposed to be the same, we can conclude that Switch 2 is chosen because it is less expensive than Switch 1, which is set to On. From P.M. 5:30 to 5:40, Switch 1 is set to On. This is considered to be to promote the snow melting at the starting point  $x=0$ [m]. Then, from P.M. 5:50 to 6:00, the switch is set to Off because the snow melting is almost completed.
We show the temporal variation of the control input  $\sigma_{{\rm pv}}(t)$ of the PV power generation equipment in Fig. 13 (b). In the time interval from P.M. 5:00 to 5:30, the equipment transmits power to the underground power distribution line. The reason for this is considered to be the increase in the power consumption of residential loads during this interval. Moreover, Fig. 13 (c) shows the temporal variation of the subscript $i$ satisfying $\sigma_{{\rm battery},i}(t)=1$, i.e., the destination of the power from the battery storage. When Switch 2 is set to On, the storage always transmits the power to the heating cable.
[…]
(a) Temporal variation of switching control input
(b) Temporal variation of the destination of PV power
(c) Temporal variation of the destination of battery power
(d) Temporal variation of the amount of electricity in battery
Figure 13: Temporal variation with respect to switches, PV, and battery storage (Case 2)
Comparison with and without battery storage
In the following, we compare the costs and the voltage fluctuations with and without battery storage. Similarly to Case 1, the former and latter simulations correspond to the authors’ previous study [9] and this paper, respectively.
[…]
Figs. 14 (c) and (d) show the voltage fluctuations without and with battery storage, respectively. From these figures, it can be confirmed that the voltage fluctuations are more suppressed when the battery storage is installed. This is because the battery storage optimizes the power flow according to the power consumption of the house and the amount of the PV power generation.
(a) Temporal variation of switching control input (without battery storage)
(b) Temporal variation of switching control input (with battery storage)
(c) Voltage fluctuation (without battery storage)
(d) Voltage fluctuation (with battery storage)
Figure 14: Comparison of switching control input and voltage fluctuation with respect to battery storage (Case 2)
V Conclusions
In this paper, we have studied the application of a battery storage to the predictive switching control of a power distribution system including road heating proposed by the authors [9]. The introduction of a battery storage increases the number of transmission destinations of power electricity from the PV generation equipment, and improves the power flow in terms of optimization. The predictive switching control is designed for the PV power generation equipment and the battery storage, respectively. Its effectiveness is demonstrated by numerical simulations using the actual time-series data. As a conclusion, we have confirmed the effectiveness of the introduction of battery storage compared to the authors’ previous study [9].
As a future work, in order to improve the control performance of the switching predictive control, it is necessary to add constraints on the voltage amplitude and to consider the weights of the evaluation function according to the distribution system and its specific situation.